OpenStudy (anonymous):

Which expression is the simplified form of sqrt 108x^3y^8?

sam (.sam.):

\[\sqrt{108 x^3 y^8}\] \[=\sqrt{108}\sqrt{x^3y^8}\] Find some gcf of 108, that is \[=\sqrt{36}\sqrt{3}\sqrt{x^3y^8}\] \[=6\sqrt{3}\sqrt{x^3y^8}\]

OpenStudy (anonymous):

ok ty